OpenStudy (anonymous):

How can i get the other values of hyperbolic functions using this tanhx=1/2 ??

OpenStudy (anonymous):

tanhx = sinhx / coshx

OpenStudy (anonymous):

= 1/2(e^x - e^-x) / 1/2(e^x + e^-x) = (e^x - e^-x) / (e^x + e^-x) = (e^2x - 1 / (e^2x + 1) Let e^2x = a (a-1)/(a+1) = 1/2 => a = 2 1/2 = (e^2x - 1 / (e^2x + 1) => 1/2 (e^2x + 1) = (e^2x - 1) => 1(e^2x + 1) = 2(e^2x - 1) => e^2x + 12 = 2e^2x - 13 => e^2x = 2 e^x = root2 => e^-x = 1/root 2 substitute get the answer

OpenStudy (anonymous):

How did it become (e^2x - 1 / (e^2x + 1) ?